# references.bib

### v1

@@ IV EXPERIMENTS
For each item, we sample 40 different random poses for the arm and template. For each pose the 4 cameras record one RGB-D image each for a total of 160 RGB-D images. In addition, we record the joints positions of the arm to correctly locate the mesh shape template in the modelled scene.
Training and inference uses an NVIDIA’s V100 GPU. All results presented are for models trained for 90K iterations which amounts to roughly 22 hours.
We present our results in two sections. Firstly, we evaluate the reconstruction fidelity from scene capture images and illustrate the re-posed capabilities. Secondly, we evaluate the impact of our positional embeddings and view direction augmentation through an ablation study.
@@ IV-A Rendering
The performances in image reconstruction are summed-up in table I and we present samples of training data reconstructions in Fig. 4a for each sock. The metrics used for comparison are the Peak Signal to Noise Ratio (PSNR) and the Structural Similarity Index Measure (SSIM) as they are common metrics for evaluating the fidelity of reconstructed images. Additionally, we report the geometric error to convey the quality of the reconstructed scene’s geometry. The geometric error is defined at the mean of the absolute value of the signed distance function’s value computed on surface sample points obtained from depth images. Higher values of PSNR and SSIM demonstrate a higher reconstruction quality while lower geometric error show better geometry reconstruction as surface points should have a signed distance of zero. All results are reported as the mean and standard deviation of the metric over 40 random images from the sock 0 dataset.
We observe from the low value of the error loss in table I that our models correctly capture the scene’s geometry. Thanks to the integration of data over multiple views and regularity constraints, our architecture is able to produce depth images that are less noisy than the original training images. We can also observe that our architecture is able to capture fine colour details such as the blue stripes on sock 0, the ripples on sock 2, or the red text on the ankle on sock 3. However we can see that the design with thinner lines in sock 1 has not been correctly captured as reflected by the lower PSNR and SSIM scores for this idem.
@@
Figure 5: (a) During the capture phase, the Franka-Emika Panda arm poses the shape template in random poses to expose all angles of the deformable object to the four RealSense cameras. (b) We have selected four sock with different patterns of varying size to display the performances of our model.
In Fig. 4b we present the results of rendering re-posed meshes with various amounts of rigidity in the material. The rendering of full resolution images (640×480px) takes roughly 20 minutes per image.
Although, the underlying poses generated by the simulator might seem unnatural for cotton socks, the depth images produced by GARField show a correct representation of the geometry of the re-posed socks. The distinctive colour features such as the blue stripes are correctly placed on the item. However, we can notice both in re-constructing the training images and in generating the re-posed images that there is some blurring of the visual feature of the socks and this erases the smaller details such as the size information present on the bottom of the sole.
Figure 6: Outputs of GARField. Top left: color. Top right: normals. Bottom left: depth. Bottom right: segmentation masks.
@@ IV-B Ablation study
To evaluate the impact of the use of the Laplacian embeddings we compare them with random embeddings and learned embeddings. Random embeddings are a set of vectors randomly sampled before training. Learned embeddings are initialized randomly and jointly optimised with the rendering model.
Each positional embedding is evaluated by training the same network architecture for 90K iterations. The metrics are computed on 40 random images of the training dataset for sock 0 (see Fig. 5b). We report the average value and standard deviation over the dataset in table II. These results confirm that the Laplacian-based position embeddings outperform the other options. By looking at the qualitative data in Fig. 7, we notice that the architecture completely fails to learn the correct geometry and appearance for the non-Laplacian embeddings. This can be attributed to the proximity-preserving nature of Laplacian-based embeddings that simplifies greatly the task of the MLP in learning a correct surface map.
@@
To evaluate the effectiveness of view direction augmentation, we proceed to three trainings on the sock 0 (see Fig. 5b) dataset with a probabilities of 0, 0.3, and 0.6 respectively for view direction augmentation. All other parameters are kept constant. From the quantitative results presented in table III we notice a slight degradation in performance in view reconstruction as the augmentation probability grows. However, looking at the sample qualitative result of a novel view rendering presented in Fig. 8, we can observe that the decrease in performance is much lower with higher augmentation percentage, as expected.
@@ V CONCLUSIONS
In this paper, we have presented a novel data generation pipeline for garments based on GARField, a fully differentiable neural scene model. For autonomous robots interacting freely with their environment, this tool can be seen as enabling robots to “imagine” how a manipulation plan of a newly discovered object would translate in the observation space. In a more immediate perspective, this tool allows for the creation of labelled datasets of deformed garments observations from simulated states. This is achieved mainly by exploiting the invariance of visual features in a surface-attached coordinate system.
The main limitation of our work is the computational heaviness of the scene rendering and scene capture. We believe that better sampling strategies could be developed to mitigate these challenges.
@@
Figure 8: Comparison an details of novel view rendering with different amount of view direction augmentation during training. Top: Full image rendering. Bottom: Colour and surface normals detail

### v2

@@ IV EXPERIMENTS
We sample 40 random poses for each item for the arm and template. For each pose, the four cameras record one RGB-D image each for a total of 160 RGB-D images. In addition, we recorded the joint positions of the arm to locate the mesh shape template correctly in the modelled scene.
Training and inference use one NVIDIA’s V100 GPU. All results are obtained after training for 90K iterations, which amounts to roughly 22 hours.
We present our results in two sections. Firstly, we evaluate the reconstruction fidelity from scene capture images and illustrate the re-posed capabilities. Secondly, we assess the effect of our positional embeddings and view direction augmentation through an ablation study.
@@ IV-A Rendering
The performances in image reconstruction are summed up in table I, and we present samples of training data reconstructions in Fig. 4a for each sock. The metrics used for comparison are the Peak Signal to Noise Ratio (PSNR) and the Structural Similarity Index Measure (SSIM), as they are common metrics for evaluating the fidelity of reconstructed images. Additionally, we report the geometric error to convey the quality of the reconstructed scene’s geometry. The geometric error is defined as the mean of the absolute value of the signed distance function’s value computed on surface sample points obtained from depth images. Higher values of PSNR and SSIM demonstrate a higher reconstruction quality, while a lower geometric error shows better geometry reconstruction, as surface points should have a signed distance of zero. All results are reported as the mean and standard deviation of the metric over 40 random images from the sock 0 dataset.
We observe from the low value of the error loss in table I that our models correctly capture the scene’s geometry. Thanks to the integration of data over multiple views and regularity constraints, our architecture is able to produce depth images that are less noisy than the original training images. We can also observe that our architecture is able to capture fine colour details such as the blue stripes on sock 0, the ripples on sock 2, or the red text on the ankle on sock 3. However, we can see that the design with thinner lines in sock 1 has not been correctly captured, as reflected by the lower PSNR and SSIM scores for this item.
@@
Figure 5: (a) During the capture phase, the Franka-Emika Panda arm poses the shape template in random poses to expose all angles of the deformable object to the four RealSense cameras. (b) We have selected four socks with different patterns of varying sizes to display the performances of our model.
In Fig. 4b, we present the results of rendering re-posed meshes with various amounts of rigidity in the material. The rendering of full-resolution images (640×480px) takes roughly 20 minutes per image.
Although the underlying poses generated by the simulator might seem unnatural for cotton socks, the depth images produced by GARField show a correct representation of the geometry of the re-posed socks. The distinctive colour features, such as the blue stripes, are correctly placed on the item. However, we can notice both in reconstructing the training images and in generating the re-posed images that there is some blurring of the visual feature of the socks, and this erases the smaller details, such as the size information present on the bottom of the sole.
Figure 6: Corresponding outputs of GARField. Top left: colour. Top right: normals. Bottom left: depth. Bottom right: segmentation masks. Black zone in the mask image corresponds to areas outside of the sphere of radius 1 centered on the scene.
@@ IV-B Ablation study
To evaluate the impact of using the Laplacian embeddings, we compare them with random and learned embeddings. Random embeddings are a set of vectors randomly sampled before training. Learned embeddings are initialised randomly and jointly optimised with the rendering model.
Each positional embedding is evaluated by training the same network architecture for 90K iterations. The metrics are computed on 40 random images of the training dataset for sock 0 (see Fig. 5b). We report the average value and standard deviation over the dataset in table II. These results confirm that the Laplacian-based position embeddings outperform the other options. Looking at the qualitative data in Fig. 7, we notice that the architecture fails to learn the correct geometry and appearance for the non-Laplacian embeddings. This can be attributed to the proximity-preserving nature of Laplacian-based embeddings, which significantly simplifies the task of the MLP in learning a correct surface map.
@@
To evaluate the effectiveness of view direction augmentation, we proceed to three trainings on the sock 0 (see Fig. 5b) dataset with probabilities of 0, 0.3, and 0.6, respectively, for view direction augmentation. All other parameters are kept constant. From the quantitative results presented in table III, we notice a slight degradation in performance in view reconstruction as the augmentation probability grows. However, looking at the sample qualitative result of a novel view rendering presented in Fig. 8, we can observe that the decrease in performance is much lower with a higher augmentation percentage, as expected.
@@ V CONCLUSIONS
This paper presents a novel data generation pipeline for garments based on GARField, a fully differentiable neural scene model. For autonomous robots interacting freely with their environment, this tool can enable robots to “imagine” how a manipulation plan of a newly discovered object would translate in the observation space. From a more immediate perspective, this tool allows for creating labelled datasets of deformed garments observations from simulated states. This is achieved mainly by exploiting the invariance of visual features in a surface-attached coordinate system.
The main limitation of our work is the computational heaviness of the scene rendering and scene capture. Better sampling strategies could be developed to mitigate these challenges.
@@
Figure 8: Comparison of details of novel view rendering with different amounts of view direction augmentation during training. Top: Full image rendering. Bottom: Colour and surface normals detail
@@ ACKNOWLEDGMENT
The authors gratefully acknowledge the financial support from the Italian Ministero degli Affari Esteri e della Cooperazione Internazionale (MAECI) as part of the project LEARN-ASSIST (PGR07303).